Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis

Road Gripper Tire Co. manufactures automobile tires. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 4,160 tires were as follows:

Standard Costs Actual Costs
Direct materials 100,000 lbs. at $6.40 101,000 lbs. at $6.50
Direct labor 2,080 hrs. at $15.75 2,000 hrs. at $15.40
Factory overhead Rates per direct labor hr.,
based on 100% of normal
capacity of 2,000 direct
labor hrs.:
Variable cost, $4.00 $8,200 variable cost
Fixed cost, $6.00 $12,000 fixed cost

Each tire requires 0.5 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Price variance $
Quantity variance $
Total direct materials cost variance $

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Rate variance $
Time variance $
Total direct labor cost variance $

c. Determine the variable factory overhead controllable variance, fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $
Fixed factory overhead volume variance $
Total factory overhead cost variance $
Answer #1

Answer:

a)

Material Price Variance = (SP-AP)*AQ purchased
MPV = (6.40-6.5)*101000 = $10100 (Unfavorable)

Material Quantity Variance = (SQ for actual output - AQ) SP
MQV = (100000-101000)*6.40 = $6400 (Unfavorable)

Total Direct Material cost variance = 10100+6400 = $16500 (Unfavorable)

b)

Labor Rate Variance = (SR-AR)AH
LRV = (15.75-15.40)2000 = -700 (Favorable)

Labor Efficiency Variance = (SH-AH)SR
LEV = (2080-2000)15.75 = -$1260 (Favorable)

Total Direct Labor cost variance = 700+1260 = -$1960 (Favorable)

c)

Variable Over Head Controllable Variance = Actual Variable OH - (SH*SR)
8200-(0.50*4*4160) = -120 (Favorable)

Fixed Volume Variance =(Capacity hours - Actual Hours)*SR
(2000-2000)*6 = 0

Total Over Head Variance = Variable + Fixed
120+0 = -120 (Favorable)